# Benchmarking Federated Learning for Semantic Datasets: Federated Scene Graph Generation

### v1

@@ V-C Convergence Behavior
We present the convergence behaviors of four models on the shard and Dirichlet distribution-based partition method in Fig. 6. When we compare the non-IID and IID cases of shard, GPS-Net shows remarkable performance improvement in IID. GPS-Net has three key modules (DMP, NPS, ARM). In prior work [47], DMP was the key module for the high performance in VG150, which has the direction of predicates in the dataset (e.g., of in hair-of-man, has in man-has-head). Because these predicates are removed in PSG dataset, the DMP module has a lower effect on performance. However, we conjecture that other modules (NPS, ARM) designed to solve long-tailed problems are effective in the FL scenarios. As a result, the performance of GPS-Net shows the fastest convergence behavior. Despite these modules, GPS-Net shows a decrease in convergence speed under the non-IID situation. MOTIFS and VCTree show similar behaviors in IID and non-IID cases, and they also have the same model structures(LSTM) and do not consider the long-tailed problem. These two methods do not seem to be significantly affected in terms of convergence speed by the non-IID situation. IMP also showed a critical decrease in convergence speed under the non-IID situation.
@@
V-D Various FL Scenarios
TABLE III: Comparison of performances for the number of total clients.
@@ V-D Various FL Scenarios
Total clients: We evaluate performances to examine the effect of number of total clients, i.e., 50, 100, and 200, in Table III. For the case of 50 clients, each client has twice as much data per client compared to the 100 clients case. Performance improves as the number of data samples that the clients have increased. Notably, VCTree shows a 2×2$\times$2 × larger mR@K for 50 clients compared to 200 clients in the Shard non-IID case, indicating that VCTree is highly sensitive to the number of data samples, a critical factor in FL settings. In contrast, MOTIFS and GPS-Net are less affected by the number of data samples, with GPS-Net achieving an mR@K greater than 5.09 and MOTIFS exceeding 4.65, significantly outperforming both IMP and VCTree across all cases.
@@
TABLE IV: Comparison of performances for participation rates
@@
We also evaluate the performance according to the number of participating clients, i.e., 5 and 20, in Table IV. As the number of participants increased, the performance of MOTIFS improved remarkably. According to previous FL studies, mainly handled the image classification task, increasing participation rate leads to improved performance in FL environments. However, in the PSG task, as the number of users increases, the performance does not show similar behaviors without MOTIFS. In other words, rather than increasing the number of participants in each round, a larger amount of data for each participant can result in greater performance improvement in this task.
@@
V-E Analysis of Communication Cost
TABLE V: Comparison of the Communication Cost for Heterogeneity
  Method # of model parameters Shard IID Shard non-IID
  IMP 32M 64(x 1) 64(x 1)
MOTIFS 63M 63(x 0.98) 63(x 0.98)
VCTree 59M 59(x 0.92) 59(x 0.92)
GPS-Net 37M 37(x 0.57) 37(x 0.57)
From the convergence behavior in Fig. 6 and 7, we compute the required communication costs to compare the communication efficiency of each method. Specifically, we calculate the total communication cost, i.e., the number of model parameters multiplied by the communication round, to reach the ‘mR@100 = 2’ (the reason for the target performance is that the IMP shows the worst performance and converges near 2). We show the number of model parameters for each method and the resulting communication costs required in Table V.
V-F Cluster Imbalance Effect
TABLE VI: Comparison of the performances on the Imbalanced Dataset
  R/mR @K Method Shard
CL[47] Random IID non-IID
     R/mR @20 IMP 16.5 / 6.52 16.10 / 5.68 16.38 / 5.97 15.50 / 4.75
MOTIFS 20.0 / 9.10 16.66 / 6.52 16.64 / 6.60 16.34 / 6.32
VCTree 20.6 / 9.70 16.49 / 6.42 16.93 / 7.03 16.46 / 6.09
GPS-Net 17.8 / 7.03 17.90 / 7.29 18.12 / 7.38 17.98 / 8.10
     R/mR @50 IMP 18.2 / 7.05 17.53 / 6.10 17.74 / 6.43 16.89 / 5.11
MOTIFS 21.7 / 9.57 18.26 / 7.03 18.38 / 7.12 17.89 / 6.70
VCTree 22.1 / 10.2 17.94 / 6.84 18.47 / 7.44 17.97 / 6.52
GPS-Net 19.6 / 7.49 19.44 / 7.77 19.78 / 7.85 19.36 / 8.46
     R/mR @100 IMP 18.6 / 7.23 18.09 / 6.28 18.20 / 6.58 17.26 / 5.22
MOTIFS 22.0 / 9.69 18.75 / 7.23 18.88 / 7.23 18.39 / 6.83
VCTree 22.5 / 10.2 18.52 / 7.00 18.92 / 7.56 18.47 / 6.64
GPS-Net 20.1 / 7.67 19.89 / 7.86 20.14 / 7.94 19.83 / 8.62
TABLE VII: Comparison of the FedAvg and FedAvgM performances of PSG methods.
  R/mR@K Method FedAvg FedAvgM[17] FedAdam[20]
Shard non-IID Shard non-IID Shard non-IID
R/mR@20 IMP 11.26 / 2.28 13.23 / 3.83 (+1.55%) 13.32 / 4.78 (+2.50%)
MOTIFS 13.33 / 4.06 15.47 / 5.80 (+1.74%) 15.89 / 5.56 (+1.50%)
VCTree 12.49 / 3.99 15.39 / 5.66 (+1.67%) 15.53 / 5.09 (+1.10%)
GPS-Net 14.57 / 5.90 16.18 / 5.91 (+0.01%) 15.66 / 5.98 (+0.08%)
R/mR@50 IMP 12.57 / 2.59 14.73 / 4.24 (+1.65%) 15.03 / 5.41 (+2.82%)
MOTIFS 14.92 / 4.48 17.23 / 6.23 (+1.75%) 17.66 / 6.01 (+1.53%)
VCTree 14.04 / 4.41 17.02 / 6.10 (+1.69%) 16.95 / 5.40 (+0.99%)
GPS-Net 16.37 / 6.36 18.00 / 6.33 (-0.03%) 17.31 / 6.42 (+0.06%)
R/mR@100 IMP 13.06 / 2.68 15.32 / 4.38 (+1.70%) 15.63 / 5.57 (+2.89%)
MOTIFS 15.43 / 4.65 17.83 / 6.38 (+1.73%) 18.21 / 6.14 (+1.49%)
VCTree 14.62 / 4.54 17.58 / 6.30 (+1.76%) 17.42 / 5.53 (+0.99%)
GPS-Net 16.91 / 6.49 18.68 / 6.52 (-0.03%) 17.90 / 6.55 (+0.06%)
()($\cdot$)( ⋅ ) indicates the difference in mR@K when each algorithm is applied compared to FedAvg.
Benchmark setups: To observe the effects of cluster imbalance, we do not equalize the number of data points. We tested 3 types of data partitioning as follows:
Results: The results in Table VI differ in some aspects from the analysis in the main paper. Firstly, the performance of all methods in FL scenarios has significantly increased. This is because the quantity of data assigned to each client has greatly increased, and models are overfitted to the dominant cluster. The performance trends of each method have changed as follows: In the case of IMP, the performance gap between CL and IID has significantly narrowed. While the performance in FL scenarios has greatly improved, the performance of CL has remained unchanged. This indicates that IMP is relatively less affected by the amount of data. In the case of MOTIFS and VCTree, the performance trends were almost similar to those observed in the previous experiments on the balanced dataset. VCTree still appears to be slightly more vulnerable to data heterogeneity compared to MOTIFS. These methods, i.e., IMP, MOTIFS, and VCTree, show that performance decreases as data heterogeneity increases. In contrast, the trends observed for GPS-Net were completely different from what was expected generally in FL scenarios. The performance reported in the previous study [47] is lower than the performance on the balanced dataset. Prior research mentioned that the key point of GPS-Net explicitly models the direction of predicates, which is why it does not perform well on PSG dataset. However, when examining the results in our main table, modeling the direction of predicates might not be the cause. Furthermore, GPS-Net showed the best results in the non-IID scenario. This is unusual and significantly deviated from our expectations. Therefore, we concluded that GPS-Net performs well when clients have sufficient data and a relatively small number of categories. Additionally, we believe that a detailed performance comparison through FL on the SGG task, rather than PSG task, would allow for a more in-depth analysis.
VI Extension to FL algorithms
From Table II, we confirm that the improvements in PSG methods remain effective in the FL scenario. Subsequently, it is necessary to verify whether the improvements in FL algorithms are also valid within our benchmark. We conducted additional experiments on two FL algorithms that leverage momentum. Momentum-based update strategies are expected to be effective in maintaining local training closer to the global update direction. By mitigating the adverse effects of data distribution discrepancies, these algorithms can potentially enhance both convergence stability and model performance in such challenging settings.
VI-A FedAvgM
We present the result of applying FedAvgM [17] in Table VII. FedAvgM utilizes the momentum in updating a global model on the server side and relieves the varying directions of local updates due to the stochastic variance across clients. FedAvgM updates the global model as follows:
$w_{g}^{r+1}=w_{g}^{r}-v^{r},$ (6)
$v^{r}=\beta v^{r-1}+\sum_{k=1}^{K}\frac{n_{k}}{n}\Delta w_{k}^{r}$ (7)
Results: FedAvgM sufficiently improves the performance of all methods. For R/mR@20, R/mR@50, and R/mR@100, there are average performance improvements of +1.24%, +1.27%, and +1.30% in the Shard-nonIID case, respectively. These performance improvements under FedAvgM can be attributed to FedAvgM’s momentum-based updates which help stabilize training by reducing local update oscillations in the optimization process, leading to faster convergence and better generalization across heterogeneous client data. IMP, MOTIFS, and VCTree showed noticeable performance increases, while GPS-Net did not. The performance of GPS-Net in the Shard non-IID case shows a negligible gap (i.e., $\leq 0.03\%$), indicating that GPS-Net already incorporates factors that mitigate the effects of heterogeneity.
VI-B FedAdam
FedOpt[20] provides a flexible framework for improving optimization in FL, enabling the use of advanced server-side optimization algorithms (e.g., AdaGrad, Adam) to improve convergence and stability. This approach effectively deals with various client data distribution differences and fluctuations in client participation rates. FedOpt uses different optimizers in local updates and global updates. In our case, we utilize the Adam[67] optimizer for global updates:
$m^{r+1}$ $=\beta_{1}m^{r}+(1-\beta_{1})\Delta w^{r},$ (8)
$v^{r+1}$ $=\beta_{2}v^{r}+(1-\beta_{2})(\Delta w^{r})^{2},$ (9)
$w_{g}^{r+1}$ $=w_{g}^{r}-\eta\frac{m^{r+1}}{\sqrt{v^{r+1}}+\epsilon},$ (10)
where $\beta_{1}$ and $\beta_{2}$ are the momentum hyperparameters, $\epsilon$ is a small constant added to the denominator to ensure numerical stability and prevent division by 00. We present the result of applying FedAdam in Table VII.
Results: FedAdam demonstrated a marginally superior performance improvement compared to FedAvgM. For R/mR@20, R/mR@50, and R/mR@100, there are average performance improvements of +1.30%, +1.35%, and +1.36% in the Shard-nonIID case, respectively. Interestingly, FedAdam shows a noticeable performance improvement when combined with IMP, as shown in the table. IMP has a lower initial performance (based on R/mR@20, R/mR@50, and R/mR@100) when compared to other methods (MOTIFS, VCTree, GPS-Net). However, when combined with FedAdam, it showed the greatest performance improvement (+2.89% in R/mR@100 case). IMP simply learns by iteratively updating relationships between objects. As a result, IMP is prone to learning by being overly head-class-biased in class imbalances, and performance degradation is inevitable in tail classes. In this environment, FedAdam has most likely improved its model effectively in the tail class, where losses are concentrated due to the long-tailed problem. Contrary to IMP, GPS-Net has various strategies to solve the long-tailed problem. Similar to the FedAvgM experimental result, GPS-Net showed no significant change in performance.
From these results, we can conclude that the enhancement of FL algorithms is effective when dealing with scenarios involving diverse semantic information across clients. Furthermore, the experimental outcomes with IMP and GPS-Net reveal an intriguing connection: the long-tailed problem encountered in scene graph generation tasks shares notable similarities with the data heterogeneity issues faced in FL. In scenarios where scene graph generation tasks must be addressed in a distributed data environment, selecting a scene graph generation method with a strong strategy for handling long-tailed problems or choosing an FL algorithm that effectively deals with data heterogeneity would significantly increase the likelihood of simultaneously tackling both challenges.
VII Discussion: Why FL Benchmarks for Scene Understanding is Required
@@
VIII Conclusion
This study introduces an innovative benchmark for evaluating FL algorithms on complex semantic datasets in the computer vision domain, particularly in the scene graph generation tasks. The proposed benchmark leverages semantic-based data clustering and controlled data partitioning to systematically manage semantic heterogeneity across clients, providing the first-ever FL benchmark for the PSG task.
The experimental results demonstrate that the benchmark effectively simulates the data heterogeneity, revealing critical insights into the performance of various PSG and FL algorithms under different scenarios. Notably, GPS-Net showcased remarkable robustness in FL settings, while simpler methods like IMP suffered significant performance degradation due to semantic heterogeneity. These findings underline the importance of addressing data heterogeneity in vision tasks and highlight the role of FL algorithms in enabling robust and decentralized training for complex tasks.
Furthermore, experiments with advanced FL algorithms, such as FedAvgM and FedAdam, revealed their effectiveness in mitigating data heterogeneity issues and improving model performance in FL scenarios. FedAdam, in particular, excelled at addressing the challenges posed by long-tailed data distributions, yielding substantial improvements for simpler methods like IMP. Conversely, GPS-Net exhibited minimal performance gains with these FL algorithms, reflecting its inherent robustness against data heterogeneity in design. These results emphasize the importance of selecting FL and PSG methods with complementary strengths to handle both semantic complexity and data heterogeneity effectively.
Acknowledgments
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2021-0-00852, Development of Intelligent Media Attributes Extraction and Sharing Technology), (No. RS2020-II201336, Artificial Intelligence Graduate School Program (UNIST)), and (No. IITP-2024-RS-2022-00156361, Innovative Human Resource Development for Local Intellectualization program), and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00459023).

### v2

@@ C.1 Convergence Behavior
We present the convergence behaviors of four models on the shard and Dirichlet distribution-based partition method in Fig. 5. When we compare the non-IID and IID cases of shard, GPS-Net shows remarkable performance improvement in IID. GPS-Net has three key modules (DMP, NPS, ARM). In prior work [28], DMP was the key module for the high performance in VG150, which has the direction of predicates in the dataset (e.g., of in hair-of-man, has in man-has-head). Because these predicates are removed in the PSG dataset, the DMP module has a lower effect on performance. However, we conjecture that other modules (NPS, ARM) designed to solve long-tailed problems are effective in the FL scenarios. As a result, the performance of GPS-Net shows the fastest convergence behavior. Despite these modules, GPS-Net shows a decrease in convergence speed under the non-IID situation. MOTIFS and VCTree show similar behaviors in IID and non-IID cases, and they also have the same model structures(LSTM) and do not consider the long-tailed problem. These two methods do not seem to be significantly affected in terms of convergence speed by the non-IID situation. IMP also showed a critical decrease in convergence speed under the non-IID situation.
@@
C.2 Analysis of Communication Cost
Table 5: Comparison of the Communication Cost for Heterogeneity
  Method # of model parameters Shard IID Shard non-IID
  IMP 32M 64(x 1) 64(x 1)
MOTIFS 63M 63(x 0.98) 63(x 0.98)
VCTree 59M 59(x 0.92) 59(x 0.92)
GPS-Net 37M 37(x 0.57) 37(x 0.57)
From the convergence behavior in Fig. 5 and 6, we compute the required communication costs to compare the communication efficiency of each method. Specifically, we calculate the total communication cost, i.e., the number of model parameters multiplied by the communication round, to reach the ‘mR@100 = 2’ (the reason for the target performance is that the IMP shows the worst performance and converges near 2). We show the number of model parameters for each method and the resulting communication costs required in Table 5.
C.3 Cluster Imbalance Effect
Table 6: Comparison of the performances on the Imbalanced Dataset
  R/mR @K Method Shard
CL[28] Random IID non-IID
     R/mR @20 IMP 16.5 / 6.52 16.10 / 5.68 16.38 / 5.97 15.50 / 4.75
MOTIFS 20.0 / 9.10 16.66 / 6.52 16.64 / 6.60 16.34 / 6.32
VCTree 20.6 / 9.70 16.49 / 6.42 16.93 / 7.03 16.46 / 6.09
GPS-Net 17.8 / 7.03 17.90 / 7.29 18.12 / 7.38 17.98 / 8.10
     R/mR @50 IMP 18.2 / 7.05 17.53 / 6.10 17.74 / 6.43 16.89 / 5.11
MOTIFS 21.7 / 9.57 18.26 / 7.03 18.38 / 7.12 17.89 / 6.70
VCTree 22.1 / 10.2 17.94 / 6.84 18.47 / 7.44 17.97 / 6.52
GPS-Net 19.6 / 7.49 19.44 / 7.77 19.78 / 7.85 19.36 / 8.46
     R/mR @100 IMP 18.6 / 7.23 18.09 / 6.28 18.20 / 6.58 17.26 / 5.22
MOTIFS 22.0 / 9.69 18.75 / 7.23 18.88 / 7.23 18.39 / 6.83
VCTree 22.5 / 10.2 18.52 / 7.00 18.92 / 7.56 18.47 / 6.64
GPS-Net 20.1 / 7.67 19.89 / 7.86 20.14 / 7.94 19.83 / 8.62
Benchmark setups: To observe the effects of cluster imbalance, we do not equalize the number of data points. The rationale for cluster balancing is that the imbalance across clusters probably prevents rigorous evaluations of FL’s capability to handle semantic heterogeneity because a model becomes overfitted to dominant clusters without balanced training across different semantics. Consequently, we believe that the balancing step is required to get rid of the overfitting issue. Also, we want to emphasize that the balanced setting, even with a quite smaller size, our result in 6 show sufficient and acceptable training behavior, i.e., shows regular convergence behavior and the discriminated final performances.
We tested 3 types of data partitioning as follows:
Results: The results in Table 6 differ in some aspects from the analysis in the main paper. Firstly, the performance of all methods in FL scenarios has significantly increased. This is because the quantity of data assigned to each client has greatly increased, and models are overfitted to the dominant cluster. The performance trends of each method have changed as follows: In the case of IMP, the performance gap between CL and IID has significantly narrowed. While the performance in FL scenarios has greatly improved, the performance of CL has remained unchanged. This indicates that IMP is relatively less affected by the amount of data. In the case of MOTIFS and VCTree, the performance trends were almost similar to those observed in the previous experiments on the balanced dataset. VCTree still appears to be slightly more vulnerable to data heterogeneity compared to MOTIFS. These methods, i.e., IMP, MOTIFS, and VCTree, show that performance decreases as data heterogeneity increases. In contrast, the trends observed for GPS-Net were completely different from what was expected generally in FL scenarios. The performance reported in the previous study [28] is lower than the performance on the balanced dataset. Prior research mentioned that the key point of GPS-Net explicitly models the direction of predicates, which is why it does not perform well on PSG dataset. However, when examining the results in our main table, modeling the direction of predicates might not be the cause. Furthermore, GPS-Net showed the best results in the non-IID scenario. This is unusual and significantly deviated from our expectations. Therefore, we concluded that GPS-Net performs well when clients have sufficient data and a relatively small number of categories. Additionally, we believe that a detailed performance comparison through FL on the SGG task, rather than PSG task, would allow for a more in-depth analysis.
C.4 Various FL Scenarios
Table 7: Comparison of performances for the number of total clients.
@@ C.4 Various FL Scenarios
Total clients: We evaluate performances to examine the effect of number of total clients, i.e., 50, 100, and 200, in Table 7. For the case of 50 clients, each client has twice as much data per client compared to the 100 clients case. Performance improves as the number of data samples that the clients have increased. Notably, VCTree shows a $2\times$ larger mR@K for 50 clients compared to 200 clients in the Shard non-IID case, indicating that VCTree is highly sensitive to the number of data samples, a critical factor in FL settings. In contrast, MOTIFS and GPS-Net are less affected by the number of data samples, with GPS-Net achieving an mR@K greater than 5.09 and MOTIFS exceeding 4.65, significantly outperforming both IMP and VCTree across all cases.
@@
Table 8: Comparison of performances for participation rates
@@
We also evaluate the performance according to the number of participating clients, i.e., 5 and 20, in Table 8. As the number of participants increased, the performance of MOTIFS improved remarkably. According to previous FL studies, mainly handled the image classification task, increasing participation rate leads to improved performance in FL environments. However, in the PSG task, as the number of users increases, the performance does not show similar behaviors without MOTIFS. In other words, rather than increasing the number of participants in each round, a larger amount of data for each participant can result in greater performance improvement in this task.
@@
C.5 Ablation Study on Number of Clusters
Table 9: Ablation studies on the number of clusters of the proposed FL benchmark.
  R/mR @K Method n = 3 n = 5 n = 10
IID non-IID IID non-IID IID non-IID
     R/mR @20 IMP 14.34 / 4.26 10.91 / 2.03 12.62 / 3.20 11.26 / 2.28 10.23 / 1.93 11.00 / 2.38
MOTIFS 14.41 / 4.72 13.90 / 3.85 13.26 / 4.64 13.33 / 4.06 11.02 / 3.07 10.82 / 2.98
VCTree 14.08 / 4.73 14.37 / 4.72 13.00 / 4.57 12.49 / 3.99 11.17 / 3.21 11.29 / 3.02
GPS-Net 16.43 / 6.71 15.83 / 6.05 14.83 / 6.90 14.57 / 5.90 12.85 / 4.93 12.55 / 4.28
     R/mR @50 IMP 15.77 / 4.65 12.30 / 2.33 13.97 / 3.53 12.57 / 2.59 11.53 / 2.18 12.36 / 2.66
MOTIFS 15.97 / 5.14 15.48 / 4.24 14.82 / 5.06 14.92 / 4.48 12.45 / 3.41 12.23 / 3.39
VCTree 15.56 / 5.09 15.90 / 5.05 14.50 / 4.94 14.04 / 4.41 12.68 / 3.61 12.63 / 3.37
GPS-Net 17.85 / 7.08 17.36 / 6.41 16.42 / 7.37 16.37 / 6.36 14.29 / 5.28 14.04 / 4.70
     R/mR @100 IMP 16.27 / 4.76 12.83 / 2.44 14.45 / 3.65 13.06 / 2.68 12.05 / 2.30 12.90 / 2.78
MOTIFS 16.48 / 5.26 16.03 / 4.40 15.38 / 5.20 15.43 / 4.65 13.02 / 3.58 12.92 / 3.58
VCTree 16.01 / 5.20 16.37 / 5.20 14.97 / 5.05 14.62 / 4.54 13.21 / 3.77 13.22 / 3.54
GPS-Net 18.31 / 7.19 17.80 / 6.51 17.08 / 7.55 16.91 / 6.49 14.85 / 5.41 14.66 / 4.83
We have repeated to test our partitioning pipeline on the PSG dataset using different number of clusters, e.g., $n=3$ and $n=10$ semantic clusters, and conducted the federated learning. We present the results on Table 9. In both cases the qualitative ordering of PSG algorithms—GPS-Net >> MOTIFS $\approx$ VCTree >> IMP—remains unchanged, and the IID $\leftrightarrow$ non-IID performance gap scales smoothly with the strength of the induced heterogeneity. However, very fine clustering ($n=10$) exaggerates data-scarcity effects and masks the heterogeneity gap. These outcomes confirm that, regardless of the exact number of clusters used and or how sharply they are separated, our pipeline continues to provide a valid data splits of federated learning with a semantic heterogeneity.
Appendix D Discussion: Why FL Benchmarks for Scene Understanding is Required
@@
Figure 7: Semantic clusters of CelebA dataset
Appendix E Details on CelebA Experiments
E.1 Clustered CelebA Dataset
We also applied our clustering method to the CelebA dataset, which is composed of 40 attributes (e.g., Eyeglasses, Wearing hat, Wavy hair, etc) and does not have a specific label for each image. Therefore, we build a non-IID dataset with the following steps. First, we perform K-Means clustering with 5 clusters. Since the attributes consist of binary values (-1, 1), we do not apply super-classes to preserve the meaning of each attribute. Second, we cluster images and allocate the cluster index to each image. Based on the cluster index of each image, we build non-IID cases by applying data partitioning methods such as shard-based and Dirichlet distribution-based partitions. Figure 7 illustrates the PCA clustering result. The qualitative visualization clearly demonstrates that our K-means Clustering leads to the splits given the semantic information.
E.2 CelebA in FL
We evaluate the classification task to check whether each person is smiling in Table 3. Similar to the result of the PSG task, it shows degradation of performance and a slow convergence rate in the shard non-IID case. In prior works, the CelebA dataset is distributed only according to identity in a non-IID case, e.g., each client has pictures of the same person, which limits the method for representing a kind of non-IID setting. Additionally, it can’t represent non-IID cases, such as CCTV, or the behavior of small devices, because all identities of each client are the same. Our clustering method builds a more practical environment, not a fixed one. For the hyperparameter, we set the local learning rate as $0.01$, momentum ratio for FedAvgM as $0.7$, and global learning rate as $1e^{-4}$ for FedAdam in $1e^{-1},1e^{-2},1e^{-3},1e^{-4}]$. The total number of rounds is $100$, and the participation ratio is $5\%$ in total number of clients $100$.